# Probing the polarized emission from the accretion-powered pulsar 4U 1907+09 with IXPE

### v1

@@ 3.4 Energy-dependent polarimetric properties in different pulse phases
We have observed energy-dependent polarimetric properties in the phase-averaged analysis and in Sect. 3.3, we concluded that they are unlikely to be produced by flares. It is then necessary to examine the energy-dependent polarimetric properties in each pulse phase. Accordingly, we divide the phase-tagged photon events into eight phase bins, as in Table 6, and into five energy bins covering the 2–8 keV range, following Tables 4 and 5. The pcube algorithm is then used to compute the PA and PD in each energy-phase bin. The results are shown in Fig. 8.
Due to the limited photon counts, most PDs in the energy-phase bins do not exceed the minimum detectable polarization (MDP) at the 99% confidence level (). Nonetheless, the evolution of PA with energy is still observable in each phase bin. The most notable detection of a PA rotation between adjacent energy bands within a single phase bin occurs in the phase interval 0.25–0.375. In this interval, the PA of photons in the 4–5 keV and 5–6 keV energy bands differs by approximately 90°. Their PDs marginally reach the , suggesting a “probable” right-angle rotation of the PA between these energy bands within this phase interval. Protractor plots illustrating the PAs with their 1, 2, and 3 uncertainties are shown in Fig. 9, where the PAs are compatible only within the 3 range. It is worth noting that in Vela X-1, a PA shift of between two energy bands is also detected within a single phase bin (Forsblom et al., 2025).
@@ 4 Discussion
XRBs are considered prime candidates for X-ray polarimetric analysis, as the highly magnetized matter accreting onto the magnetic poles of the NS can naturally generate a high PD due to the birefringence of the medium in such a strong magnetic field. The oscillations of the electric field of the emitted photons are expected to align either parallel or perpendicular to the plane formed by the photon momentum and the local magnetic field (Gnedin & Pavlov, 1974; Gnedin et al., 1978). These modes are referred to as the ordinary (O) and extraordinary (X), respectively. The Compton scattering cross-section below the cyclotron energy for the O-mode radiation is much lower than that for the X-mode (Lai & Ho, 2003) predicting a high PD up to 80% (Meszaros et al., 1988; Caiazzo & Heyl, 2021).
Nevertheless, most XRBs observed by IXPE exhibit low phase-averaged PDs (see, e.g., Tsygankov et al., 2022; Mushtukov et al., 2023; Forsblom et al., 2023), with higher PDs emerging in phase-resolved analyses. This trend does not fully hold for 4U 1907+09, where polarization is weakly detected in most phase bins. Further examination of the energy dependence of the PA within individual phase bins partially enhances the significance of polarization detection in some energy-phase bins, as shown in Fig. 8. However, some energy-phase bins still exhibit extremely low PD significance.
@@
Various accretion processes involving reflection and scattering can introduce additional polarized components that tangle with the emission produced near the magnetic poles (see the discussion in Tsygankov et al., 2022). Consequently, it is natural to consider an energy-dependent polarized component as a potential explanation for the unusual energy dependence of the PA. This idea has been explored by Forsblom et al. (2025) and Loktev et al. (2025). Specifically, Forsblom et al. (2025) proposed two possible mechanisms for the observed PA variation with energy. The first involves a spectral component that diminishes at higher energies and generally has a PA orthogonal to that of the dominant high-energy component. The second invokes vacuum resonance, in which one polarization mode converts into another near the cyclotron line energy, resulting in a shift in PA. Given the results shown in Tables 4 and 5, it is difficult to envision a spectral component that is extraordinarily strong specifically in the 3–4 and 5–6 keV bands, thereby causing the observed drop in PDs. Regarding the second mechanism, it predicts an energy-independent PA at both low and high energies. However, this prediction cannot be confirmed due to the limited photon statistics ( million photons within the 2–8 keV band for the combined dataset) and the restricted energy coverage of IXPE. Furthermore, if a gradual shift in PA with energy is observed, it would indicate the presence of an additional polarized component. More robust results are expected from future observations with improved photon statistics and broader energy coverage.
@@ 5 Summary
4U 1907+09 was observed by IXPE two times on 2024 November 11–21 during its brightest periods close to the periastron. We summarize the results of the polarimetric analysis of 4U 1907+09 as follows:
@@
A more significant polarization is detected in Obs. 1 compared to Obs. 2, with phase-averaged PD and PA values of and for Obs. 1, and and for Obs. 2. The combined data set from both observations yields a PD of and a PA of .
@@
Energy-dependent PA is likely observed in the phase-averaged polarimetric analyses with a significance of 1.7 , partially explaining the generally insignificant polarization detections. Specifically, we detect a probable rotation of the PA between adjacent energy bands within one single phase bin.
@@
We further examine whether the three brief flares detected by IXPE significantly affect the pulsar’s polarization. Our analysis concludes that the polarimetric properties are similar in both flare-excluded and flare-only data, suggesting that the observed energy-dependent PAs are unlikely to be attributed to the flares.
@@ Acknowledgements.
- Arnaud (1996) Arnaud, K. A. 1996, in ASP Conf. Ser., Vol. 101, Astronomical Data Analysis Software and Systems V, ed. G. H. Jacoby & J. Barnes (San Francisco: Astron. Soc. Pac.), 17–20
@@
- Caiazzo & Heyl (2021) Caiazzo, I. & Heyl, J. 2021, MNRAS, 501, 109
@@
- Elsner & Lamb (1977) Elsner, R. F. & Lamb, F. K. 1977, ApJ, 215, 897
@@
- Gnedin & Pavlov (1974) Gnedin, Y. N. & Pavlov, G. G. 1974, Soviet Journal of Experimental and Theoretical Physics, 38, 903
@@
- Marshall & Ricketts (1980) Marshall, N. & Ricketts, M. J. 1980, MNRAS, 193, 7P
@@
- Mushtukov & Tsygankov (2024) Mushtukov, A. & Tsygankov, S. 2024, Accreting strongly magnetised neutron stars: X-ray Pulsars, ed. C. Bambi & A. Santangelo (Singapore: Springer), 4105–4176
@@
- Naghizadeh-Khouei & Clarke (1993) Naghizadeh-Khouei, J. & Clarke, D. 1993, A&A, 274, 968
@@
- Radhakrishnan & Cooke (1969) Radhakrishnan, V. & Cooke, D. J. 1969, Astrophys. Lett., 3, 225

### v2

@@ 3.4 Energy-dependent polarimetric properties in different pulse phases
We have observed energy-dependent polarimetric properties in the phase-averaged analysis, and in Sect. 3.3 we concluded that they are unlikely to be produced by flares. It is then necessary to examine the energy-dependent polarimetric properties in each pulse phase. Accordingly, we divided the phase-tagged photon events into eight phase bins, as in Table 6, and into five energy bins covering the 2–8 keV range, following Tables 4 and 5. The pcube algorithm was then used to compute the PA and PD in each energy-phase bin. The results are shown in Fig. 8.
Due to the limited photon counts, most PDs in the energy-phase bins do not exceed the minimum detectable polarization at the 99% confidence level (). Nonetheless, the evolution of the PA with energy is still observable in each phase bin. The most notable detection of a PA rotation between adjacent energy bands within a single phase bin occurs in the phase interval 0.25–0.375. In this interval, the PA of photons in the 4–5 keV and 5–6 keV energy bands differs by approximately 90°. Their PDs marginally reach the , suggesting a “probable” right-angle rotation of the PA between these energy bands within this phase interval. Protractor plots illustrating the PAs with their 1, 2, and 3 uncertainties are shown in Fig. 9, where the PAs are compatible only within the 3 range. It is worth noting that in Vela X-1, a PA shift of between two energy bands is also detected within a single phase bin (Forsblom et al., 2025).
@@ 4 Discussion
X-ray Pulsars are considered prime candidates for X-ray polarimetric analysis, as the highly magnetized matter accreting onto the magnetic poles of the NS can naturally generate a high PD due to the birefringence of the medium in such a strong magnetic field. The oscillations of the electric field of the emitted photons are expected to align either parallel or perpendicular to the plane formed by the photon momentum and the local magnetic field (Gnedin & Pavlov, 1974; Gnedin et al., 1978). These modes are referred to as the ordinary (O) and extraordinary (X), respectively. The Compton scattering cross section below the cyclotron energy for the O-mode radiation is much higher than that for the X-mode (Lai & Ho, 2003), predicting a high PD up to 80% (Meszaros et al., 1988; Caiazzo & Heyl, 2021).
Nevertheless, most XRPs observed by IXPE exhibit low phase-averaged PDs (see, e.g., Tsygankov et al., 2022; Mushtukov et al., 2023; Forsblom et al., 2023), with higher PDs emerging in phase-resolved analyses. This trend does not fully hold for 4U 1907+09, where polarization is weakly detected in most phase bins. Further examination of the energy dependence of the PA within individual phase bins partially enhances the significance of polarization detection in some energy-phase bins, as shown in Fig. 8. However, some energy-phase bins still exhibit extremely low PD significance.
@@
Various accretion processes involving reflection and scattering can introduce additional polarized components that tangle with the emission produced near the magnetic poles (see the discussion in Tsygankov et al., 2022). Consequently, it is natural to consider an energy-dependent polarized component as a potential explanation for the unusual energy dependence of the PA. This idea has been explored by Forsblom et al. (2025) and Loktev et al. (2025). Specifically, Forsblom et al. (2025) propose two possible mechanisms for the observed PA variation with energy. The first involves a spectral component that diminishes at higher energies and generally has a PA orthogonal to that of the dominant high-energy component. The second invokes vacuum resonance, in which one polarization mode converts into another near the cyclotron line energy, resulting in a shift in the PA. Given the results shown in Tables 4 and 5, it is difficult to envision a spectral component that is extraordinarily strong specifically in the 3–4 and 5–6 keV bands, thereby causing the observed drop in PDs. Regarding the second mechanism, it predicts an energy-independent PA at both low and high energies. However, this prediction cannot be confirmed due to the limited photon statistics ( million photons within the 2–8 keV band for the combined dataset) and the restricted energy coverage of IXPE. Furthermore, if a gradual shift in the PA with energy is observed, it would indicate the presence of an additional polarized component. More robust results are expected from future observations with improved photon statistics and broader energy coverage.
@@ 5 Summary
4U 1907+09 was observed by IXPE twice during the period 2024 November 11–21, during its brightest periods close to the periastron. We summarize the results of the polarimetric analysis of 4U 1907+09 as follows:
@@
A more significant polarization is detected in Obs. 1 than in Obs. 2, with phase-averaged PD and PA values of and for Obs. 1 and and for Obs. 2. The combined dataset from the two observations yields a PD of and a PA of .
@@
An energy-dependent PA is likely observed in the phase-averaged polarimetric analyses with a significance of 1.7 , partially explaining the generally insignificant polarization detections. Specifically, we detect a probable rotation of the PA between adjacent energy bands within one single phase bin.
@@
We further examined whether the three brief flares detected by IXPE significantly affect the pulsar’s polarization. Our analysis concludes that the polarimetric properties are similar in the flare-excluded and flare-only data, suggesting that the observed energy-dependent PAs are unlikely due to the flares.
@@ Acknowledgements.
- Arnaud (1996) Arnaud, K. A. 1996, in ASP Conf. Ser., Vol. 101, Astronomical Data Analysis Software and Systems V, ed. G. H. Jacoby & J. Barnes (San Francisco: Astron. Soc. Pac.), 17
@@
- Caiazzo & Heyl (2021) Caiazzo, I., & Heyl, J. 2021, MNRAS, 501, 109
@@
- Elsner & Lamb (1977) Elsner, R. F., & Lamb, F. K. 1977, ApJ, 215, 897
@@
- Gnedin & Pavlov (1974) Gnedin, Y. N., & Pavlov, G. G. 1974, Soviet Journal of Experimental and Theoretical Physics, 38, 903
@@
- Marshall & Ricketts (1980) Marshall, N., & Ricketts, M. J. 1980, MNRAS, 193, 7P
@@
- Mushtukov & Tsygankov (2024) Mushtukov, A., & Tsygankov, S. 2024, Accreting strongly magnetised neutron stars: X-ray Pulsars, ed. C. Bambi & A. Santangelo (Singapore: Springer), 4105
@@
- Naghizadeh-Khouei & Clarke (1993) Naghizadeh-Khouei, J., & Clarke, D. 1993, A&A, 274, 968
@@
- Radhakrishnan & Cooke (1969) Radhakrishnan, V., & Cooke, D. J. 1969, Astrophys. Lett., 3, 225